### v3

# Breaking the Trend: How to Avoid Cherry-Picked Signals

Abstract
Our empirical results show an impressive fit with the pretty complex theoretical Sharpe formula of a Trend following strategy depending on the parameter of the signal, which was derived by Grebenkov and Serror (2014). That empirical fit convinces us that a mean-reversion process with only one time scale is enough to model, in a pretty precise way, the reality of the trend-following mechanism at the average scale of CTAs and as a consequence, using only one simple $\mathrm{EMA}$, appears optimal to capture the trend. As a consequence, using a complex basket of different complex indicators as signal, do not seem to be so rational or optimal and exposes to the risk of cherry-picking.
August 7, 2025
I thank Professors Campbell Harvey, Robert Kosowsky, Valeriy Zakamulin, Denis Grebenkov and Raphael Douady for their feedbacks and valuable insights.
1 Introduction
Practitioners of the trend-following strategies usually optimize their strategies empirically by selecting parameters that maximize the simulated reward-to-risk ratio, typically measured by the Sharpe ratio, when replaying the historical scenarii through past returns. However, this empirical approach can lead to overfitting, especially when optimization is performed on the same historical dataset used for evaluation.
An alternative is to model the returns of the underlying assets using stochastic processes. In some cases, this allows the Sharpe ratio to be derived mathematicaly through a formula depending on the parameters of the strategy. When this is possible, we refer to it as the theoretical Sharpe ratio, in contrast to the empirical one.
It is also possible to determine the set of parameters that maximizes the theoretical Sharpe ratio—this is known as the theoretical optimal Sharpe ratio, as opposed to the common empirical optimum obtained through backtesting.
A theoretical Sharpe ratio is always interesting if the model deriving the returns is realistic enough, as it allows validating and optimizing a strategy without the risk of overfiting.
The main objective of this paper is to empirically validate the theoretical Sharpe ratio model proposed by Grebenkov and Serror (2014) and therefore their assumption to model trends as gaussian mean reversion stochastic processes with only one time scale.
2 Literature
The usual recipe to determine a portfolio in the CTA’s industry is to use a blend of many different technical indicators as a signal for each underlying instrument (most often individual indicator gives either a long or a short position111 Sepp and Lucic (2025) classified the methodologies for sizing positions based on signals into three categories: (1) European trend following, which assumes continuous position sizes proportional to signal strength — an approach used by many large European CTA managers; (2) Time Series Momentum (Moskowitz et al (2012); Hurst et al (2013); Baltas (2015)), a method more commonly used in academic studies, where positions depend on the sign of the momentum; and (3) American trend following, which originates from the early adopters of CTA strategies.). Then, a risk management process is applied to the signals to ensure that the portfolio is diversified enough and target s a constant risk. One simple solution is to impose an equal conditional risk in each asset class, as the SG Trend Indicator does and to size the portfolio to target a constant volatility. Another solution is to use the correlation matrix and apply mathematical optimization to maximize the reward/risk ratio. The well-known Markowitz solution set positions as linearly depending on the signals. The linear dependence is derived simply througth the normalization of the signals by the inverse of the correlation matrix between returns, but the results are not appealing. The Agnostic Risk Parity ($\mathrm{ARP}$), introduced by Benichou at al (2017), normalizes the signals through the inverse of the square root of the correlation matrix. Benichou at al (2017) proposed this portfolio because it was rotationally invariant. Since this concept is primarily understood by physicists, their portfolio is not yet widely accepted in the finance community, with only a few citations, despite posing a serious challenge to Markowitz.
In reality, one hidden assumption of Markowitz optimization is that the expected returns from the signals are known and certain, but this assumption is incorrect. Valeyre (2024) proved that the Agnostic Risk Parity ($\mathrm{ARP}$) approach was optimal when the correlation matrix between the trends or signals was a kind of random matrix where a dominant factor of this correlation matrix was very dominant and diffused randomly. Valeyre (2024) suggested an interacting agents model to justify such an hypothesis.
Most of the popular technical indicators are based on moving averages of past prices. The most popular is the Simple Moving Average ($\mathrm{SMA}$), while less commonly used types include the Linear Moving Average and the Exponential Moving Average on returns ($\mathrm{EMA}$). Each moving average is computed using an averaging window of a particular size. Trend-following indicators can also be based on a combination of moving averages, such as crossovers (one with a short window size and another with a long window size). The switch is determined when the short-term moving average crosses the long-term moving average. As an example, The Societe Generale Trend signal, which is a reference among CTAs, is detemined only by the crossover based on a simple average based 20 and 120 day parameters. Momentum ($MOM$) is also a very popular indicator and is defined as the difference between two prices. Hurst et al (2017) analysed the performance simulated on the last century using a mixture of 1 month, 3 month and 1 year momentum. Lamperiere et al (2014) analysed the performance simulated on the last 2 centuries using a 50 day EMA signal. The Bollinger Band ($\mathrm{BB}$) is very popular, but non continuous with only 3 outputs 0 and 1 or -1 when price is oustide a band. The Moving Average Convergence/Divergence ($\mathrm{MACD}$) uses a combination of three EMAs to capture mean reversion at both short and long-term scales while identifying trends in the long term.
We can describe some of these indicators through their sensitivity to past daily returns as introduced by Zakamuli and Giner (2020). Fig.1 displays some cases. It can be observed that the Exponential Moving Average ($\mathrm{EMA}$) exhibits a sensitivity that declines exponentially, while the crossover resembles a hat shape. Sensitivity is low for very recent past returns, high for a moderately distant past, and negligible for a long past. The shape of the crossover can be adapted if price behavior exhibits mean-reverting tendencies on very short time scales, as seen in the case of individual stocks with low liquidity. From an academic point of view, almost all these trend-following rules are ad hoc and lack theoretical justification for their optimality. Zakamuli and Giner (2020) related them to each other through the analysis of their sensitivity to past daily returns (so a 33 day EMA should be the optimal EMA to replicate the mixuture of 1 month, 3 month and 1 year momentum used by Hurst et al (2017) while a 43 day EMA should be the optimal one for the SG trend Indicator crossover). Zakamuli and Giner (2020) also reminds us that it is well known that if returns follow an autoregressive process, the best predictor has the same functional form as the autoregressive process being predicted. Therefore, the expected structure of the autoregressive process should provide a good proxy for the best predictor.
Zakamuli and Giner (2024) shows that the indicator with the same sensitivity to past returns as the autoregressive coefficients of the return process is both the best predictor and the optimal indicator, yielding the highest Sharpe ratio when the investor has only two options (Buy or Sell with the same level of risk). He also employs a two-state regime-switching model (bull and bear regimes), a widely accepted framework for stock returns (Timmermann (2000); Fruhwirth-Schnatter (2006); Giner and Zakamulin (2023)), using a semi-Markov model to detect negative autocorrelations over very long time scales. He argues that, within this framework, the MACD indicator with three different time scales could be adapted to construct the optimal portfolio, but obtaining statistically significant results remains difficult. So mean reversion at larger scales remains unconvincing.
However, the two-state regime does not seem natural for modeling trend-following mechanisms, which are better explained by herding behavior rather than by an external guiding force. Models in which agents partially imitate each other are more likely to produce continuous regime changes rather than abrupt shifts between two states. Consequently, a more realistic approach would require a more complex model than a simple two-state regime. A more credible representation would assume that trends follow an Ornstein-Uhlenbeck process or a similar mean-reverting process.
Grebenkov and Serror (2014) proved that, in the more complex yet more insightful case where portfolio positions depend linearly on the signal, the $\mathrm{EMA}$ with a larger decay parameter than the decay parameter of the autoregressive coefficients of the return process yields the optimal portfolio. This finding is in stark disagreement with Zakamuli and Giner (2024), simply because Grebenkov and Serror (2014) solved a different, yet superior, optimization problem. Indeed, assuming that positions depend linearly on the signal appears much more natural and is not as suboptimal as merely assuming a binary choice between a long or short position with the same level of risk withtout the possibility of implementing “money management”. Most important the version of Grebenkov and Serror (2014) is in line with Markowitz optimization and Agnostic risk parity by Benichou at al (2017) where positions are linearly depending on the signals.
Moreover, Grebenkov and Serror (2014) is the only one to have derived an elegant and rigorous formula for the theoretical Sharpe ratio, explicitly linking it to the parameters of the return process and the relaxation time of the $\mathrm{EMA}$. That formula is key in our main objective as we want to validate it using empirical measurements.
Two years later, while Grebenkov and Serror (2014) remained largely unrecognized by the academic finance community, the literature continued to focus on the wrong optimization problem—one where positions do not depend linearly on the signals. Dai at al (2016) addressed this issue by imposing a constraint that allowed only two possible positions (long or zero). Similarly, Nguyen at al (2014b) determined the optimal threshold for triggering either a long or short position.
Our first objective was to validate empirically the model of Grebenkov and Serror (2014) through testing their beautiful formula describing the sensitivity of the empirical Sharpe ratio of the “Agnostic Risk Portfolio” ($\mathrm{ARP}$)—one of the components of the optimal trend-following portfolio derived by Valeyre (2024)—to the parameter of the $\mathrm{EMA}$. We also wanted to explore more complex signals than $\mathrm{EMA}$ to confirm that one time scale is enough to describe trends. These empirical results challenge the common belief in our industry that complex systems are necessary. For example, Tzotchev (2018) derived an optimization scheme for assigning weights across different time scales using a HRP-Markowitz framework, based on theoretical correlations results derived from an AR(1) process for returns. It also challenges the very controversial  [^2] Buncic (2025); Elder (2025) empirical and theoretical findings of Kelly et al (2023), who theoretically argue that simple models severely understate return predictability. We decided to limit the other signals to a combination of three $\mathrm{EMA}$s to align with the $\mathrm{MACD}$ which could make sense if the autoregressive structure of returns could be more complex than the one which fits perfectly for a $\mathrm{EMA}$. However, there are two major differences:
  • We aim to set the slope of the sensitivity at lag zero to zero as we focus only on systematic returns through cross asset futures excluding specific risk. We prefer to avoid a positive slope, as it would be curious for recent returns to have less impact on the trend than older ones and as short term Mean reversion behavior is known to be less important than for single stocks.
  • We aim to achieve a larger fat tail for the larger scales instead of having a contrarian contribution which appears more adapted at capturing residual risk in single stocks but not for systematic part of returns.
Our second objective was to understand that a simple Exponential Moving Average ($\mathrm{EMA}$), which is supposed to be both optimal and simple, could also be replicated by a highly complex but more usual combination of indicators, such as Bollinger Bands, whose elementary indicator is even nonlinear and path-dependent. This second objective raises the question: why use a complex combination of complex indicators that are sensitive to cherry-picking when a simple and elegant solution already exists?
(a) MOM 100 days
(b) SMA 100 days
(c) EMA 100 days
(d) SMA Crossover 30-100 days
(e) EMA Crossover 30-100 days
(f) MACD 30-100-400 days
Figure 1: Capture of Zakamuli and Giner (2024). “Spectrumm” of different indicators based on Zakamuli and Giner (2024) i.e. sensistivity to past daily returns. $\mathrm{MOM}$ is the different between two prices, $\mathrm{SMA}$ is the usual moving average on prices, $\mathrm{EMA}$ is the exponential moving avarage of returns, $\mathrm{SMA}$ crossover is the difference between 2 $\mathrm{SMA}$, $\mathrm{EMA}$ crossover is the difference between 2 $\mathrm{EMA}$. $\mathrm{MACD}$ has 3 time scales and could be adapted according Zakamuli and Giner (2024) which is contrarian on the large scales
3 Theory from Grebenkov and Serror (2014)
To derive the theoretical Sharpe ratio for a trend-following strategy, we need a model that describes returns and trends. A classical one-dimensional model used to capture trends in an underlying asset assumes that returns consist of two components: a random, independent and identically distributed (i.i.d.) component, and a small bias term. This bias can be modeled using an auto-regressive process of order 1, AR(1), which represents mean reversion process around zero with only one time scale. The AR(1) process is characterized by one time scale parameter—indicating the average time needed to revert to zero—and a standard deviation that determines the level of noise. AR(1) is very similar if time lag is infinitesimally small to the Orstein-Uhlenbeck process which is a mean reversion usual and natural process in complex systems with only one time scale.
This classical one-dimensional model can be easily extended to multiple dimensions. In the multivariate case, two covariance matrices are required to account for correlations. The first matrix captures the correlations between the random components of the different underlyings, while the second matrix reflects the correlations among the AR(1) processes associated with each underlying. The model is formally described in Eq. 1, using the same notation as in Grebenkov and Serror (2014). In that equation the returns $r_{i,t}$ of the $i$th instrument among the N is described in two components: The first $\epsilon_{i,t}$ is a noise which is not autocorrelated while the second $\beta\sum_{k=1}^{t-1}{\left(1-\lambda\right)^{t-1-k}\xi_{i,k}}$ is an AR(1) process which is autocorrelated where $\xi_{i,k}$ is a noise which is not autocorrelated. Ultimately Eq.1 describes the returns of a diffusive process with a positive autocorrelation model for returns $r_{i,t}$: The weights $\left(1-\lambda\right)^{t-1-k}$ are declining exponentialy. The parameter $\lambda$ describes the inverse of the relaxation time of the mean reversion AR(1) process describing the short bias, and $\beta$ represents the strength of that short stochastic bias in the returns. The noises $\xi_{i,k}$ are independent in the $k$ axis but may be correlated with $i$ axis (Valeyre (2024) introduces specific forms of that matrix, leading to particular optimal portfolios). The noises $\epsilon_{i,t}$ are independent in the $t$ axis but may be correlated with $i$ axis with a very close correlation matrix between returns. Through that model, the trend is $\beta\sum_{k=1}^{t-1}{\left(1-\lambda\right)^{t-1-k}\xi_{i,k}}$ and it follows an auto-regressive model of order 1, AR(1).
$r_{i,t}=\epsilon_{i,t}+\beta\sum_{k=1}^{t-1}{\left(1-\lambda\right)^{t-1-k}\xi_{i,k}}$ (1)
Note that an alternative model is also used in the literature, for example in Tzotchev (2018); Sepp and Lucic (2025), where the returns, rather than the trends, are modeled directly using an AR(1) process or even an autoregressive fractal process. However, this approach is less common than the classical model used by Grebenkov and Serror (2014).
Through variogram measurements on the Dow Jones Index, they estimated that $\lambda=0.01$ and $\beta_{0}=0.1$ (Fig.2). The fit appears suspiciously perfect, considering the expectation of multiple time scales with long memory among investors as well as the potential presence of mean-reversion patterns at longer horizons that are not captured. Additionally, the result is not entirely convincing, as the 100-year period may be heterogeneous, and the fit is less accurate at shorter time scales, where measurements should, in principle, be less noisy. Another drawback is that the fit may be specific to the Dow Jones Index, the only index with such a long history. To say it in another way, we can suspect the presence of two distinct time scales, as the fit is not perfect for time scales shorter than 100 days. Another possible explanation is that the autocorrelation may not have been consistent over such a long period.
Figure 2: Capture of Grebenkov and Serror (2014). Variogram of standardized logarithmic daily returns (normalized by realized volatility) to Dow Jones (1900-2012) fit with $\beta_{0}=0.08$ and $\lambda=0.011$ parameters
Grebenkov and Serror (2014) solved the correct theoretical optimization problem—maximizing the Sharpe ratio at the portfolio level while ensuring that positions remain linearly dependent on signals when assuming that returns are well described by Eq.1. They considered only the one-dimensional case, but later, Grebenkov and Serror (2015) and Valeyre (2024) extended the analysis to the multi-dimensional setting.
Instead of focusing on the best predictor or on the optimal portfolio with suboptimal constraints—such as the possibility of being either long with the same risk or flat, which is unfortunately common in the literature—Grebenkov and Serror (2014) assumes that positions should better depend linearly on the signals [^3] They considered only the EMA as the signal, which was a good choice, as Zakamuli and Giner (2024) showed that the best predictor has the same functional form as the autoregressive process being predicted. That is, in the case of Eq. 1, the optimal predictor should be an EMA, which they indeed used, which is a natural approach which is in agreement with the Markowitz solution. They then determine the optimal signal that maximizes the Sharpe ratio. Thus, Grebenkov and Serror (2014) introduced substantial novelty, even though it is not yet widely considered by most academics specializing in trend-following.
In detail, they derived an explicit formula, in Eq. 3, the theoretical Sharpe ratio (${\mathbf{S}}\left(\eta\right)$), when assuming returns are generated through their autocorrelation model in Eq. 1, as a function of $\eta$ (the smoothing parameter of the trend indicator $EMA$ which is defined mathematically later in Eq.6 ) , $\lambda$ (the inverse of the time scale of the mean reversion process that describes the trends in Eq. 1) and $\beta$ (the weight of the trend component in the returns Eq. 1) or its normalized version $\beta_{0}$ while also considering the inclusion of trading costs, denoted as $\theta$. Equation Eq.3 is in fact a remarkably elegant formula. As expected, the theoretical Sharpe ratio depends on $\beta$ (or $\beta_{0}$ its normalized version), the weight of the trend in the returns, in an approximately proportional manner when $\beta_{0}^{2}$ is significantly larger than $\frac{1}{\lambda}$ and when $\eta$ remains close to $\lambda$. In that case, the Sharpe ratio from Eq.3 can be approximated by $\beta_{0}\sqrt{\frac{1}{2}}$ or $\beta_{0}\sqrt{\frac{255}{2}}$ in the annualized version of the Sharpe. Fig.3 illustrates the very complex theoretical formula Eq.3 through a simple graph that helps to understand its sensitivity to the $\mathrm{EMA}$ parameter.. In reality, Eq. 3 was derived for the one-dimensional case only. Nevertheless, we can reasonably expect the equation to hold in the multi-dimensional setting. In that case, Eq. 2 must be adjusted to redefine $\beta_{0}$. To be consistent with the observed increase of the Sharpe ratio with the size of the universe $N$, we can, under strong approximations, expect $\beta_{0}$ to be related to $\frac{\beta}{\sqrt{\lambda\left(2-\lambda\right)}}\sqrt{\frac{N}{1+\left(N-1\right)\rho^{2}}}$ where $\rho^{2}$ is the average squared correlation in the universe. This expression extends Eq. 2 using Eq. (40) from Grebenkov and Serror (2015).
Additionally they derived in Eq.4, the optimal smoothing parameter ($\eta_{\text{opt}}$) of the Exponential Moving Average ($\mathrm{EMA}$) applied to returns as a trend indicator for a trend-following strategy which is supposed to yield to the optimal Sharpe ratio. It is interesting to observe from Eq.4 that $\frac{1}{\eta_{\text{opt}}}$, the time scale of the signal $\mathrm{EMA}$ generating the optimal strategy is always shorter than $\frac{1}{\lambda}$, the time scale of the best predictor $\mathrm{EMA}$ (or the optimal indicator in the sub optimal optimization problem of Zakamuli and Giner (2024)) and that of the autocorrelated process of returns. Indeed, when $\beta_{0}$ increases (either due to a stronger trend or a more diversified universe), the time scale of the optimal $\mathrm{EMA}$ should be reduced. As a conclusion, theoretically the parameters of the signals should be adjusted when the universe is increased.
$\beta=\beta_{0}\sqrt{\lambda\left(2-\lambda\right)}$ (2)
${\mathbf{S}}\left(\eta\right)=\frac{\beta^{2}_{0}\sqrt{2\eta}-\frac{2}{\pi}\theta\sqrt{\eta}\left(\lambda+\eta\right)}{\sqrt{\left(\lambda+\eta\right)^{2}+2\beta^{2}_{0}\left(\lambda+\eta\right)}}$ (3)
$\eta_{\text{opt}}=\lambda\sqrt{1+2\frac{\beta^{2}_{0}}{\lambda}}$ (4)
Figure 3: Capture of Grebenkov and Serror (2014). Theoretical Sharpe ($\mathbf{S}$} from Eq. 3) of the trend following strategy depending on $\eta$ the parameter of the $\mathrm{EMA}$. $\beta_{0}=0.1$ and $\lambda=0.01$ for different cost of trading $\theta$
4 Description of the empirical analysis
To validate Grebenkov’s theory (Eq. 3), introduced in Sec 3 and, therefore, the diffusive process in Eq. 1 which is assumed for the underlying assets, we conducted empirical backtests of the Agnostic Risk Parity (ARP) strategy introduced by Benichou at al (2017). This strategy constructs a rotationally invariant portfolio that applies trend-following while allocating equal unconditional risk to each eigenvector of the covariance matrix. While rotational invariance is a desirable theoretical property, the equal unconditional risk allocation among eigenvectors is a more desired property for asset managers. Moreover, Valeyre (2024) demonstrated that the ARP portfolio is optimal—achieving the highest Sharpe ratio—under the assumption that the covariance of the $\xi_{i,k}$ terms in Eq. 1 follows a particular random matrix structure, which is considered realistic. Implementing the ARP portfolio is relatively straightforward and consists of normalizing the trend indicators by the inverse square root of the correlation matrix, rather than by the full inverse as in the classical Markowitz framework.
We decided to focus primarily on exponential moving averages ($\mathrm{EMA}$) as trend indicators to feed the Agnostic Risk Parity (ARP) strategy, as they are theoretically optimal according to Grebenkov and Serror (2014); Zakamuli and Giner (2024), assuming the validity of Eq. 1. Accordingly, we tested various $\mathrm{EMA}$ configurations and measured how the empirical Sharpe ratio of the resulting portfolio depends on the EMA parameter, in line with Grebenkov’s theoretical predictions. Additionally, we explored more complex trend indicators based on multiple time scales—such as the Moving Average Convergence Divergence ($\mathrm{MACD}$)—to empirically validate whether the $\mathrm{EMA}$ remains the optimal choice.
We backtested the Agnostic Risk Parity (ARP) strategy on a global universe including Futures on commodities, fx, stock indices and bonds on a period 25th of May 1990 and stops on the 7th of December 2023.
4.1 Renormalized exponential moving average indicator ($\mathrm{EMA}$)
We used only a normalized version of the $EMA$ to account for the dependence of volatility on both time and instrument. This normalization allows for meaningful comparisons of trend indicators across instruments, time periods, and even different smoothing parameters. At time $t$, the $\mathrm{EMA}$ is a vector of $N$ exponential moving averages of normalized returns, one for each of the $N$ instruments. The $EMA$ of instrument $i$ at time $t+1$ is defined by $\phi$ in Eq. 6, where $r_{i,t+1}$ denotes the return of instrument $i$ at time $t+1$. The parameter $\sigma$ represents the standard deviation of returns used to normalize them, and $\eta$ is the smoothing parameter of the $EMA$ as well as of the return normalization. $\sigma^{2}$ can be interpreted as an exponential moving average of squared returns, and the indicator $\phi$ corresponds to an exponential moving average of returns scaled by their volatility. The incrementation is on a daily basis but it could be generalized to minutes returns.
$\sigma^{2}_{i,t+1}\left(\eta\right)=\left(1-\eta\right)\sigma^{2}_{i,t+1}\left(\eta\right)+\eta r^{2}_{i,t+1}$ (5)
$\phi_{i,t+1}\left(\eta\right)=\left(\ 1-\eta\right)\phi_{i,t}\left(\eta\right)+\sqrt{\eta}\frac{r_{i,t+1}}{\sigma_{i,t}\left(\eta\right)}$ (6)
The indicator $\mathrm{EMA}\left(150\right)$ is defined by the $\phi_{i,t+1}\left(\eta\right)$ when $\eta=\frac{1}{150}$ is applied to every underlying $i$ and time $t$. We use $\sqrt{\eta}$ in Eq.6 so that the std of $\phi_{i,t+1}\left(\eta\right)$ is 1 theoretically if $r$ are well approximated by unautocorrelated returns so that $\phi$ is a normalized $EMA$ indicator.
4.2 Moving Average Convergence Divergence indicator ($\mathrm{MACD}$) as a 3 time scales $\mathrm{EMA}$ indicator
Inspired by $\mathrm{MACD}$, we introduced a combination of $\mathrm{EMA}$ applying the Eq.7 while determining $\omega_{1}$ so that the derivative of the sensitivity to past daily returns is at 0 at the lag 0 and replicate curves in Fig.4. The idea behind this indicator was to reduce the weight of returns at very short time scales—where mean reversion is possible—and to increase the weight of returns at longer time scales, in order to test a multi-timescale memory effect for trend-followers.
$\mathrm{MACD}_{i,t}\left(\eta_{1},\eta_{2},\eta_{3},\omega_{1},\omega_{2},\omega_{3}\right)=\omega_{1}\phi_{i,t}\left(\eta_{1}\right)+\omega_{2}\phi_{i,t}\left(\eta_{2}\right)+\omega_{3}\phi_{i,t}\left(\eta_{3}\right)$ (7)
The derivative at zeros yields to Eq.8.
$0=\omega_{1}\sqrt{\eta_{1}}+\omega_{2}\sqrt{\eta_{2}}+\omega_{3}\sqrt{\eta_{3}}$ (8)
4.3 Agnostic Risk Parity ($\mathrm{ARP}$) when smoothing the portfolio and targeting a constant volatility
We first estimated the correlation matrix $C$, of dimension $N\times N$, using a 750-day exponential moving average applied to weekly returns when implemnting the RIE filter introduced by Bun at al (2016). The vector $\Sigma$ consists of N values representing the standard deviations estimated using a 40-day exponential moving average applied to daily returns.
Next we expressed the positions vector as a linear function of the signals ($\mathrm{EMA}$ through $\phi$ or $\mathrm{MACD}$) applying a normalisation that involves the inverse of the square root of the estimated correlation matrix and volatilities . This follows the formulation of the agnostic risk parity portfolio ($\mathrm{ARP}$}) introduced by Benichou at al (2017).
We then used the parameter $\rho=\frac{1}{20}$ for portfolio smoothing as specified in Eq.LABEL:eq:arp2 which is an very easy solution to reduce trading cost (usually, trend indicators include rules to reduce trading costs. While the $EMA$ does not explicitly aim to avoid excessive trading, smoothing the portfolio has proven to be both simple and effective. Benichou at al (2017) recommended optimization under constraints to better manage trading costs).
Finally, we applied a resizing process in Eq.13 to target a constant volatility for the final positions $\mathrm{ARP}$} which is a vector of N weights at time $t+1$.
[…]
${\mathrm{ARP}}_{t+1}=\frac{\hat{\mathrm{ARP}}_{t+1}}{\sqrt{\hat{\mathrm{ARP}}_{t+1}^{\prime}{\Sigma_{t}}C{\Sigma_{t}}\hat{\mathrm{ARP}}_{t+1}}}$ (13)
4.4 Dataset and different simulated parameters
The simulation starts on the 25th of May 1990 and stops on the 7th of December 2023. We used daily returns from 70 futures instruments in stock indices, bonds, fx and commodities futures. The description is in the appendix A.
We tested the different indicators applying the $\mathrm{ARP}$ formula Eq.13. The different parameters are described in Tab 3 in the appendix B.
Fig.4 displays the sensitivities of these indicators to past daily returns. $\mathrm{MACD}$ enables us to put more weights on very older returns as we expected.
Figure 4: Different sensitivities to past daily returns for indicators ($\mathrm{MACD}$, $\mathrm{EMA}$) all based on 20, 80 and 400 day time scales. we have $\mathrm{ARP}\left(0\times 20,80,0.2\right)$, $\mathrm{ARP}\left(0\times 20,80,0.4\times 400\right)$, $\mathrm{ARP}\left(20,80,0.2\times 400\right)$, $\mathrm{ARP}\left(20,80,0.4\times 400\right)$$\mathrm{ARP}\left(80\right)$, $\mathrm{ARP}\left(400\right)$
5 Empirical results
Here we first interpret the simulated empirical Sharpe ratio when applying the $\mathrm{ARP}$ portfolio with the different indicators. We find interesting results which challenge traditional recipes. Indeed, and quite surprisingly, the empirical Sharpe ratio closely replicates the theoretical Sharpe ratio predicted by Grebenkov and Serror (2014), whose main result shows that a single $EMA$ indicator is already optimal. Secondly, we attempt to replicate the performance of a simple $EMA$ using a basket of more complex trend indicators. Thirdly, we incorporate the non-linear findings of Schmidhuber (2021), which could challenge our assumptions, and finally, we analyze the empirical correlations between different strategies.
5.1 Grebenkov’s model empirical validation
In Section 1 and 3, we introduced both the empirical and theoretical Sharpe ratios. As a reminder, the empirical Sharpe ratio is derived from backtesting, whereas the theoretical Sharpe ratio is obtained by modeling the underlying asset using theoretical diffusive processes. We tested Eq. 3, which describes the theoretical Sharpe ratio as a function of the parameters of the trend indicator, as derived by Grebenkov and Serror (2014).
Interestingly, the empirical simulation fits pretty well the theorical formula of the Sharpe ratio Eq.3 derived in Grebenkov and Serror (2014) with the following parameters $\lambda=1/180$ and $\beta_{0}=0.12$. The Fig.5 is very impressive and the empricial fits should validate Gebenkov’s model to describe trends. The risk of overfitting is very limited, as the formula was originally derived in 2014, with its parameters fitted solely on a long historical series of the Dow Jones Index. Using an extended dataset including commodities, forex, bonds, and other stock indices, we obtain parameter estimates that are consistent with those in Grebenkov and Serror (2014), even if they differ significantly in value. Moreover, the fit appears to be more robust than that obtained using the variogram method (Fig. 2) applied to the Dow Jones Index, as our fitting is performed on a broader set of asset classes over a more recent period. That is the main result of the paper.
As a first consequence the parameter of $112$ days ($\eta_{\text{opt}}=\frac{1}{112}$) for simple $\mathrm{EMA}$ is the optimal parameter to get the optimal Sharpe ratio when not accounting cost of trading which are very small at that trading frequency (Tab. 4 displays holding periods slower than 80 days). Lamperiere et al (2014) found a different result, with a faster optimal parameter at 50 days for the EMA, but very similar Sharpe ratios across EMAs ranging from 1 to 5 months, showing an almost flat curve with no clear optimum. The difference with our results may be explained by the choice made in Lamperiere et al (2014) to determine positions based not on the linear magnitude of the EMA signals, but solely on their sign. Additionally, they did not use the ARP portfolio construction, which involves inverting the square root of the correlation matrix—a method that explains a significant part of the Sharpe ratio, as shown in Benichou at al (2017). For these two reasons, their Sharpe ratio for the post-2000 period was measured at 0.85, lower than our 1.2, and their optimal EMA signal corresponded to a faster timescale than ours.
As a second consequence of the very good fit with the theoretical formula, modeling trend through mean-reversion process using only one relaxation time and not an multi-time scales one appears to be a good solution as the fit is more than correct. That is particulary unexpected as market is known to have a multi-time scales property: For example the relaxation of volatility is known to have multi-time scales, investors are expected to have different horizons of time and different horizons of analysis. As a consequence we can wonder whether the usual recipe to take into account a multitude of different indicators is justified to claim having the most robust and the most refined signal. When properly renormalizing the signal, the position, and the portfolio, we do not observe trends reversing over longer horizons (e.g., 12 months) that would indicate mean reversion. Instead, the Grebenkov formula still holds even at horizons of up to 1,000 days. This stands in strong contradiction to the findings of Moskowitz et al (2012), who focused on momentum (MOM) indicators over the period 1965–2009.
We can also note that $\beta_{0}=0.12$ is slightly higher than $0.08$, the parameter measured for the Dow Jones over the past 100 years by Grebenkov and Serror (2014). This difference may seem minor, but since the Sharpe ratio depends on $\beta_{0}^{2}$, it results in a Sharpe ratio that is 2.25 times higher when applying the strategy to a universe of 70 underlying assets instead of the Dow Jones (assuming the trend strength over the past 30 years was similar to that of the past 100 years). This further confirms the importance of measuring the implied autocorrelation parameter based on a strategy invested in a large universe, as it leads to more accurate and agregated estimates.
We can also note that $\lambda$ is estimated to $\frac{1}{180}$ instead of $0.011$ in the case of the Dow Jones. Our analysis is that market behavior should have changed in the last 100 years and we believe that our fit appears more robust than a simple variogram.
We can see based on the Tab 4 in appendix C that empirical Sharpe is $1.24$ for $\mathrm{ARP}\left(120\right)$ with one time scale and $1.18$ for $\mathrm{MACD}\left(20,120,0.4\times 400\right)$ with 3 time scales. $\mathrm{MACD}$ could not be justified as additional time scales do not bring significant improvement. Also Sharpe ratio is not so sensitive to the parameters arround the optimal as expected: Sharpe is $1.25$ for $\mathrm{ARP}\left(100\right)$ and $1.21$ for $\mathrm{ARP}\left(150\right)$.
Figure 5: Empirical Sharpe ratio based on the whole period 1990-2023 for trend following strategies using $\mathrm{ARP}$ (Eq.13 ) applied to $\mathrm{EMA}$ as signal v.s. $\eta$ the parameter of the $\mathrm{EMA}$ (Eq.6) and the theoretical equation Eq.3. The theoretical model is fitted with parameters $\beta_{0}=0.12$ and $\lambda=1/180$ so that $\eta_{\text{opt}}=\frac{1}{112}$ . Empirical measurements are displayed in the tab 4.
5.2 Replication of a simple $\mathrm{EMA}$ by a mixture of Bollinger Bands $\mathrm{BB}$
The Bollinger Bands indicator, $\mathrm{BB}$, applies a double heavy-side function to an $\mathrm{SMA}$ with a width $\delta$. This indicator is nonlinear and therefore complex, with sensitivities to past returns that depend on the price path. However, the usual approach is to use a signal composed of a mixture of many Bollinger Bands indicators with different parameters, making the signal less path-dependent and increasingly robust. Here, we aim to demonstrate that a simple $\mathrm{EMA}$ signal can be replicated through a complex mixture of a large number of $\mathrm{SMA}$, which can, in turn, be decomposed into a large number of $\mathrm{BB}$ Bollinger Bands (Eq.16). This explains why it is common, as shown in Fig.6 when replication the optimal $\mathrm{EMA}$ with 112 days, to display indicator weights following a bell-shaped curve centered around 200 days, while explaining to investors that the signal contains both short-term and long-term indicators.
$\mathrm{EMA}_{t}\left(\eta\right)=\frac{r_{t}+\left(1-\eta\right)r_{t-1}+\left(1-\eta\right)^{2}r_{t-2}+....+\left(1-\eta\right)^{n}r_{t-n}+....}{1+\left(1-\eta\right)+\left(1-\eta\right)^{2}+....+\left(1-\eta\right)^{n}+...}$ (14)
$\mathrm{EMA}_{t}\left(\eta\right)=\frac{...+\left[\left(1-\eta\right)^{n-1}\eta\right]\left(n-1\right)\mathrm{SMA}_{t}\left(n-1\right)+...+\left[\left(1-\eta\right)\eta\right]\mathrm{SMA}_{t}\left(1\right)}{1+\left(1-\eta\right)+\left(1-\eta\right)^{2}+....+\left(1-\eta\right)^{n}+...}$ (15)
$\mathrm{SMA}_{t}\left(n\right)=\int_{0}^{\infty}\mathrm{BB}_{t}\left(n,\delta\right)\,d\delta$ (16)
Figure 6: weigths of a mixture of $\mathrm{BB}$ Vs size of the window to replicate the $\mathrm{EMA}\left(\frac{1}{112}\right)$ which is derived from Eq.16 corrected to pick different parameters of $\mathrm{BB}$ uniformely into lag in logarithm
5.3 Non-linear findings which challenge the EMAEMA
Schmidhuber (2021) provides empirical evidence across different time scales suggesting that when trends are extreme, the optimal indicator becomes negatively dependent on the trend itself. As a consequence, strong trends appear to be relatively rare. Specifically, for a normalized trend variable $\phi$, they find that the optimal signal depends on a nonlinear transformation of the form $\phi-c\phi^{3}$, with $c=0.33$. They regress daily normalized returns on both the normalized trend and its cubic term as explanatory variables, across various underlyings and asset classes, including commodities.
While they use a different trend indicator than $\phi$ derived in Eq. 6, our version of the renormalized $\mathrm{EMA}$, their results challenge the assumption of a purely gaussian diffusive return process from Eq. 1, which underpins the theoretical optimality of $\mathrm{EMA}$-based strategies. Interestingly, when applying the Agnostic Risk Parity (ARP) strategy using $\phi-c\phi^{3}$ instead of $\phi$ as the trend signal, we observe no improvement in the Sharpe ratio. This might be due to the lack of significance in the measurement of the coefficient $c$ (the sensitivity to the cubic term), which is highly sensitive to a few extreme trends, or to the loss of the rotational invariance property of the ARP (applying $\phi-c\phi^{3}$ to the raw returns or to their projections on the eigenvectors of the covariance matrix no longer yields equivalent results due to the nonlinearity of the transformation).
A classical approach to introduce mean-reverting behavior from fundamentalist traders is to assume an attraction toward a fair or efficient price though such a price is unobservable. It can be approximated by a moving average, which effectively induces negative autocorrelation, especially over longer time scales. Zakamuli and Giner (2024) explore this idea, though obtaining statistically significant results remains difficult. By contrast, Schmidhuber (2021)’s approach involves making autoregressive coefficients trend-dependent, enabling a dynamic transition between trend-following governed by trend-followers and mean-reversion regimes governed by fundamentalist traders. Another way to model the influence of fundamental investors could be simply to assume their impact manifests in changes to the trend, i.e., on the noise $\xi_{i,k}$ in Eq. 1. It could be enough to make sure that prices are not totally disconnected from fair and efficient ones which would differ in that case from moving average of prices.
The findings of Schmidhuber (2021) should not be ignored. A deeper investigation is, however, warranted to better understand their implications for the construction of optimal trend-following portfolios, although this lies beyond the scope of the present simple work and should be the subject of a dedicated study.
5.4 Correlations between indicators
Fig.7 in Appendix C shows that the indicators with different parameters give very correlated strategies which confims that the traditional approach of using a basket of many indicators as signal is not so appealing. $\mathrm{ARP}\left(80\right)$ is correlated to $\mathrm{ARP}\left(150\right)$ with a coefficient of $0.96$. That is nevertheless pretty interesting to see such strong correlations which could be explained by a common factor somewhere, which could be the object of an additional research. Morevever the $\mathrm{ARP}\left(120\right)$ based on a simple $\mathrm{EMA}$ is very close with correlation from 1 to 0.99 to the new indicator $\mathrm{MACD}\left(20,120,0\times 400\right)$ or $\mathrm{MACD}\left(20,120,0.4\times 400\right)$ we introduced based on 3 different time scales to increase weights for very old past returns and decrease weights on very recent ones. So it makes this refinement not justified again.
6 Conclusion
Grebenkov’s model for describing trends is empirically validated, as its pretty complicated theoretical formula for determining the Sharpe ratio based on the $\mathrm{EMA}$ parameter fits impressively well with empirical data. The best fit is obtained using the theoretical model parameters $\lambda=1/180$ and $\beta_{0}=0.12$. As a consequence, the parameter of $112$ days for simple $\mathrm{EMA}$ is the optimal parameter to get the optimal Sharpe ratio. It is quite surprising that a single $\mathrm{EMA}$ is optimal for capturing trends, as one would expect different time scales for different types of investors. However, there are likely much shorter time scales, on the order of a few days, but they have no significant impact on a medium-frequency strategy. The conclusion is that using a complex mixture of sophisticated indicators is unnecessary when the $\mathrm{EMA}$ alone provides a perfect fit—proving that simplicity can indeed be beautiful.
7 Declaration of funding
No funding was received
8 Disclosure of interest
There are no interests to declare,
Appendix A Data
Commodities
Brent Crude (IFEU \$/bbl)
Cocoa (IFUS \$/mt)
Coffee (IFUS \$/lbs)
Corn (CBT \$/bu)
Cotton #2 (IFUS \$/lbs)
Crude Oil WTI (NYM \$/bbl)
ECX EUA (IFEU EUR/t)
Feeder Cattle (CME \$/lbs)
Gasoil (IFEU \$/mt)
Gold (NYM \$/ozt)
Hard Red Wtr Wheat (CBT \$/bu)
High Grade Copper (NYM \$/lbs)
Iron Ore 62% Fe, CFR China (TSI) (NYM \$/mt)
Lean Hogs (CME \$/lbs)
Live Cattle (CME \$/lbs)
Lumber (CME \$/bft)
Milling Wheat (LIF EUR/t)
Natural Gas (NYM \$/mmbtu)
NY Harb RBOB (NYM \$/gal)
NY Harbor ULSD (NYM \$/gal)
Oats (CBT \$/bu)
Orange Juice (IFUS \$/lbs)
Palladium (NYM \$/ozt)
Platinum (NYM \$/ozt)
Rough Rice (CBT \$/cwt)
Rubber RSS3 (TKT JPY/kg)
Silver (NYM \$/ozt)
Soybean Meal (CBT \$/t)
Soybean Oil (CBT \$/lbs)
Soybeans (CBT \$/bu)
Sugar #11 (IFUS \$/lbs)
Wheat (Chicago) - Contract
Table 1: List of the 32 commodities futures traded in the USA or in Europe.
Stock indices Bond indices FOREX
AMSTERDAM EOE Idx
S&P Midcap 400 Idx e-mini
Russell 2000 Idx e-mini
Cac 40
Dax
Ftse 100
STOXX Europe 600 Index Futures
Hang Sen
Mib S&p-mif
Nikkei 225 Osaka
Topix
Kospi 200
Ibex 35
Mini MSCI Emerging Markets Index Future
Nasdaq E-mini
S&P 500 e-mini
Dj Euro Stoxx
S&P Canada 60-ME
SPI 200 Idx
Mini Dow Futures
BUND 10Yr
CAD Bond 10Yr
Bobl
Schatz
Long-term Euro-btp
Euro-buxl Futures
LONG Gilt 10Yr
10yr Fr Gov Bond
US T-NOTE 5Yr
JGB 10Yr
US T-NOTE 10Yr
US T-Note 2Yr
Ultra T-Bonds Combined
AUD/USD Fut.
GBP/USD Fut.
CAD/USD Fut.
EUR/USD Fut.
JPY/USD Fut.
MXN/USD Fut.
NZD/USD Fut.
CHF/USD Fut.
Table 2: List of the 41 instruments among currencies, equity indices and Bonds.
Appendix B Parameters
portfolio comment indicator type
$\mathrm{ARP}\left(20\right)$ for 20 days as relaxation time for the $\mathrm{EMA}$ when applying $\mathrm{EMA}\left(\frac{1}{20}\right)$
$\mathrm{ARP}\left(50\right)$ when applying $\mathrm{EMA}\left(\frac{1}{50}\right)$
$\mathrm{ARP}\left(80\right)$ when applying $\mathrm{EMA}\left(\frac{1}{80}\right)$
$\mathrm{ARP}\left(100\right)$ when applying $\mathrm{EMA}\left(\frac{1}{100}\right)$
$\mathrm{ARP}\left(120\right)$ when applying $\mathrm{EMA}\left(\frac{1}{120}\right)$
$\mathrm{ARP}\left(150\right)$ when applying $\mathrm{EMA}\left(\frac{1}{150}\right)$
$\mathrm{ARP}\left(180\right)$ when applying $\mathrm{EMA}\left(\frac{1}{180}\right)$
$\mathrm{ARP}\left(400\right)$ when applying $\mathrm{EMA}\left(\frac{1}{400}\right)$
$\mathrm{ARP}\left(1000\right)$ when applying $\mathrm{EMA}\left(\frac{1}{1000}\right)$
$\mathrm{ARP}\left(20,120,0\times 400\right)$ when applying $\mathrm{MACD}\left(\frac{1}{20},\frac{1}{120},\frac{1}{400},\omega_{1},1,0\right)$
$\mathrm{ARP}\left(0\times 20,120,0\times 400\right)$ when applying $\mathrm{MACD}\left(\frac{1}{20},\frac{1}{120},\frac{1}{400},0,1,0\right)$
$\mathrm{ARP}\left(20,120,0.2\times 400\right)$ when applyingr $\mathrm{MACD}\left(\frac{1}{20},\frac{1}{120},\frac{1}{400},\omega_{1},1,0.2\right)$
$\mathrm{ARP}\left(20,120,0.4\times 400\right)$ when applying $\mathrm{MACD}\left(\frac{1}{20},\frac{1}{120},\frac{1}{400},\omega_{1},1,0.4\right)$
$\mathrm{ARP}\left(20,90,0.3\times 400\right)$ when applying $\mathrm{MACD}\left(\frac{1}{20},\frac{1}{90},\frac{1}{400},\omega_{1},1,0.3\right)$
$\mathrm{ARP}\left(20,80,0.3\times 400\right)$ when applying $\mathrm{MACD}\left(\frac{1}{20},\frac{1}{80},\frac{1}{400},\omega_{1},1,0.3\right)$
$\mathrm{ARP}\left(20,80,0.2\times 400\right)$ when applying $\mathrm{MACD}\left(\frac{1}{20},\frac{1}{80},\frac{1}{400},\omega_{1},1,0.2\right)$
$\mathrm{ARP}\left(20,80,0.4\times 400\right)$ when applying $\mathrm{MACD}\left(\frac{1}{20},\frac{1}{80},\frac{1}{400},\omega_{1},1,0.4\right)$
Table 3: Different parameters fo the trend indicator
Appendix C Results
portfolio indicator type Gross Sharpe ratio average holding period (days))
$\mathrm{ARP}\left(20\right)$ $\mathrm{EMA}$ 1.079535 38
$\mathrm{ARP}\left(50\right)$ $\mathrm{EMA}$ 1.189320 60
$\mathrm{ARP}\left(80\right)$ $\mathrm{EMA}$ 1.240349 74
$\mathrm{ARP}\left(100\right)$ $\mathrm{EMA}$ 1.244945 81
$\mathrm{ARP}\left(120\right)$ $\mathrm{EMA}$ 1.235455 88
$\mathrm{ARP}\left(150\right)$ $\mathrm{EMA}$ 1.207496 96
$\mathrm{ARP}\left(180\right)$ $\mathrm{EMA}$ 1.172569
$\mathrm{ARP}\left(400\right)$ $\mathrm{EMA}$ 0.955223 132
$\mathrm{ARP}\left(1000\right)$ $\mathrm{EMA}$ 0.633678 155
$\mathrm{ARP}\left(0\times 20,120,0\times 400\right)$ $\mathrm{MACD}$ 1.235455 88
$\mathrm{ARP}\left(20,120,0.2\times 400\right)$ $\mathrm{MACD}$ 1.203418 97
$\mathrm{ARP}\left(20,120,0.4\times 400\right)$ $\mathrm{MACD}$ 1.176466
$\mathrm{ARP}\left(20,90,0.3\times 400\right)$ $\mathrm{MACD}$ 1.214172 89
$\mathrm{ARP}\left(20,80,0.3\times 400\right)$ $\mathrm{MACD}$ 1.218031 85
$\mathrm{ARP}\left(20,80,0.2\times 400\right)$ $\mathrm{MACD}$ 1.228186 82
$\mathrm{ARP}\left(20,80,0.4\times 400\right)$ $\mathrm{MACD}$ 1.206864 87
Table 4: Gross sharpe ratio based on the whole period 1990-2023
Figure 7: Empirical correlation on the whole period 1990-2023 between ARP models with different trend indicators parameters